[Lerftest] Isotope production update
jordan at jlab.org
Thu Sep 17 14:06:54 EDT 2020
Neil Wilson canceled crane due to rain tomorrow, so blocks are to stay in place till next week
Operations staff will support an additional 24 hours of operations; the LERF will be staffed until Friday end of day shift
RadCon will retrieve crucible & begin gamma spec measurements on Saturday
The 1/4 RF has been off for 12+ hours:
Quarter 3 Repair and Recovery<https://logbooks.jlab.org/entry/3850429>
Lognumber 3850429<https://logbooks.jlab.org/entry/3850429>. Submitted by kerns<https://logbooks.jlab.org/user/kerns> on Thu, 09/17/2020 - 13:32<https://logbooks.jlab.org/entries?start_date=1600360353&end_date=1600367553&book=LERFLOG&book=ELOG&book=RFLOG>.
Last updated on Thu, 09/17/2020 - 13:32
Logbooks: LERFLOG<https://logbooks.jlab.org/book/lerflog> ELOG<https://logbooks.jlab.org/book/elog> RFLOG<https://logbooks.jlab.org/book/rflog>
Entry Makers: kerns
Called in at 0700 about Quarter 3 not going into HV. Found HV enable fault not clearing and discovered Ross relay was welded closed and inoperable and preventing HV and crowbar from clearing on thyratron tube. New Ross relay found in cabinet (probably in there for 10 years !!) and installed it. No problem in going to HV except half an hour to condition thyratron to make it go into ready state. All recovered to RF and ready for B. Freeman to beam on in the LERF !
Beam is back on!!!
Lognumber 3850434<https://logbooks.jlab.org/entry/3850434>. Submitted by bfreeman<https://logbooks.jlab.org/user/bfreeman> on Thu, 09/17/2020 - 13:51<https://logbooks.jlab.org/entries?start_date=1600361479&end_date=1600368679&book=LERFLOG>.
Last updated on Thu, 09/17/2020 - 13:51
Entry Makers: bfreeman
Input for this video channel is on SLM2F09
Fig. 1 [09/17/2020 13:50:18]
Many thanks to all for the continued support & dedication to making this run a success!
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Lerftest